A divalent metal ion M2+ reacts with a unidentate ligand (L) to form a 1:1 complex ML22+. Calculate the concentration of M2+ in a solution prepared by mixing equal volumes of 0.20 M M2+ and 0.20 M L if the Kf is 1.0 x 108

1
Expert's answer
2021-07-23T08:11:01-0400

Solution


[M2+] = 0. 10 M, [L] = 0. 10 M

M2+ + L D ML2+ Kf = ( 0 [M2+] = 0.10 M, [L] = 0.10 M M2+ + L D ML2+ Kf = ( 0.10 –x )/x2


0.10>>x since kf is very large

1.0x108= 0.10/x2, x = 3.2x10-5


Relative error = "(\frac{3.2\times10^{-5}}{0.10}) \times100 = 3.2\times10^{-2}%"



[M2+] = 3.2x10-5M